# Nonlinearly charging the conformally dressed black holes preserving duality and conformal invariance

### v1

@@ Abstract
We start this paper by concisely rederiving ModMax, which is nothing but the unique nonlinear extension of Maxwell’s equations preserving conformal and duality invariance. The merit of this new derivation is its transparency and simplicity since it is based on an approach where the elusive duality invariance is manifest. In the second part, we couple the ModMax electrodynamics to Einstein gravity with a cosmological constant together with a standard conformal scalar field, and new stationary spacetimes with dyonic charges are found. These solutions are later used as seed configurations to generate nonlinearly charged (super-)renormalizably dressed spacetimes by means of a known generating method that we extend to include any nonlinear conformal electrodynamics. We end by addressing the issue of how to generalize some of these results to include the recently studied non-Noetherian conformal scalar fields, whose equation of motion still enjoys conformal symmetry even though its action does not. It results that the static non-Noetherian conformally dressed black holes also become amenable to being charged by ModMax.
@@ I Introduction
It is certainly not an understatement to say that the paradigm of all effective low-energy field theories is due to the seminal works of Heisenberg, Euler, and Weisskopf [1, 2]. In more modern language, it is agreed that by integrating out the degrees of freedom of the electron field in quantum electrodynamics (QED), one obtains an approximate description of the photon field deviating from the standard linear regime [3]. This QED vacuum behavior makes nonlinear electrodynamics a semi-phenomenological tool. As a result, several observational and experimental efforts have been deployed over the years to probe possible nonlinear effects in electrodynamics [4]. Well before these developments, Born and Infeld constructed a theory in which the electromagnetic field of a point charge is devoid of singularities [5]. Indeed, by going beyond the principle of superposition, they were able to solve the problem of infinite self-energy, and they argued that there should be a maximum field strength similar to the upper limit provided by the speed of light in special relativity. This reasoning was found to be consistent with string theory decades later, where Born-Infeld theory reappears in the dynamics of strings and branes, and this maximum field strength is precisely associated with the fact that such extended objects can not move faster than the speed of light [6, 7]. This result further strengthens the interpretation of nonlinear electrodynamics as an effective field theory.
@@
The mentioned scale dependence is incompatible with conformal symmetry, one of the key features of standard Maxwell’s theory in four dimensions and in fact, of almost the whole standard model. Hence, it is also desirable to have explicit self-gravitating configurations involving other electrodynamics that depart from the Maxwell one but still preserving the conformal invariance. It is worth mentioning that static black holes are known for electromagnetism theories with conformal symmetry for dimensions $D\not=4$, which can be achieved by considering as Lagrangian a power law of the single field strength invariant with a dimension-dependent exponent [13]. These nonlinear models have also been shown to harmonize with conformally coupled scalar fields [14], even to the point of producing composite stealth configurations. [^1] By stealth we mean, nontrivial field configurations with a vanishing energy-momentum tensor, see Refs. [15, 16, 17, 18] It is natural to consider scalar fields at the same time since they are precisely the simplest way to realize conformal symmetry. With this respect, we mention that the most general second-order conformally invariant scalar field equation in four dimensions arising from an action principle was recently found in [19] after the precursor results of Ref. [20].
@@
In the context of black holes the relevance of conformally invariant fields has been evident since the pioneering and independent works of Refs. [21, 22], where the first example of a black hole dressed by a conformal scalar field was provided. This solution turns out to be pathological due to the divergence of the scalar field at the event horizon [23]. Remarkably, this pathology can be cured just by turning on a cosmological constant together with the unique conformally invariant potential in four dimensions [24]. This spherical black hole allows an electrically charged extension [24] and an analog whose event horizon is negatively curved [25], for completeness we also mention that they have been generalized in several ways as solutions, see e.g. [26, 27, 28, 29]. Of course, the dyonic extension of this so-called charged MTZ black hole is straightforward thanks to the duality and conformal symmetries of the standard Maxwell theory. Outstandingly, it turns out that Maxwell electrodynamics is not the only one that shares such properties. In fact, there exists a unique nonlinear extension of the Maxwell theory that inherited both its conformal and duality symmetry, which was recently proposed in Ref. [30] and dubbed ModMax. The main goal of this paper is to explore if the mere assumption of these two ingredients, i.e. duality as well as conformal invariance, is enough to find further generalizations of the known conformally dressed black holes. The results of our exploration are positive and, in what follows we summarize the different steps to achieve our objective.
@@
The plan of the paper is organized as follows. In order to be self-contained but also for usefulness, we start in Sec. II by providing a survey of nonlinear electrodynamics à la Salazar, García and Plebański [31]. This step is needed since it allows a transparent definition of duality invariant theories that we will carefully justify in Subsec. III.1. The conformal symmetry is then characterized in Subsec. III.2 and later used beside the previous results to allow a clear rederivation of the ModMax theory in Subsec. III.3. In Sec. IV, we provide the dyonic generalization, supplemented with a nontrivial NUT parameter, of the MTZ black hole nonlinearly charged through the ModMax theory. Since we go beyond Maxwell theory respecting power-counting renormalizability it makes total sense to do something similar with the scalar field. Fortunately, in Ref. [28] it was shown by means of a generating technique that the charged MTZ solution allows a generalization supported by a (super-)renormalizable self-interaction potential. Here, we first extend this generating technique to include any conformal nonlinear electrodynamics and then construct new (super-)renormalizably dressed spacetimes in Sec. V. This work is finally completed by adding to the scalar field a term that breaks its conformal invariance at the level of the action but not at the level of its equation of motion, i.e. we also address the study of the so-called non-Noetherian conformal scalar fields [20, 19]. In Sec. VI we succeed to charge with ModMax the static non-Noetherian conformally dressed black holes originally found by Fernandes in Ref. [20]. The last section VII is devoted to our conclusions.
@@ II Nonlinear electrodynamics
which depends on the metric $g_{\mu\nu}$, the gauge potential $A_{\mu}$ and the antisymmetric tensor $P^{\mu\nu}$. Here the structural function $\mathscr{H}$ describes the precise nonlinear electrodynamics and depends in general on the two Lorentz scalars that can be constructed with $P^{\mu\nu}$ [5]; see the first equality of Eq. (9b). As usual, the field strength is related to the gauge potential as $F=\textrm{d}A$, ensuring the Faraday equations
@@
with the sub-indices of $\mathscr{H}$ indicating partial differentiation, e.g. ${\mathscr{H}}_{\mathscr{P}}={\partial\mathscr{H}}/{\partial\mathscr{P}}$. Notice that Maxwell electrodynamics is recovered for $\mathscr{H}=\mathscr{P}$, giving linear constitutive relations. Lastly, the corresponding energy-momentum tensor reads
@@
The main motivation for the action principle (1) is that now Maxwell equations (3) remain linear as the Faraday ones (2) while the nonlinearity is encoded into the constitutive relations (4). Consequently, Maxwell equations (3) can be now understood just like the Faraday ones (2), i.e. implying the local existence of a vector potential $\star P=\textrm{d}A^{*}$. Therefore, from the point of view of the action principle (1), a solution to nonlinear electrodynamics can be understood as a pair of vector potentials $A$ and $A^{*}$ compatible with the constitutive relations (4). Additionally, since in four dimensions both Faraday (2) and Maxwell (3) equations define conservation laws, there are conserved quantities related to them defined by the following integrals
@@
After this brief and useful introduction, and in order to prepare what follows, we review a strategy that has proved to be fruitful when nonlinear electrodynamics is considered in General Relativity [32]. This strategy simply consists of working in a null tetrad of the spacetime metric
@@
Here, the first pair of the tetrad is composed of complex conjugates one-forms, while the last pair is real. Additionally, it has been implicitly assumed that the electromagnetic configuration is algebraically general, namely, the real invariants $E$, $B$, $D$, and $H$ which are related to the eigenvalues are not all zero at the same time. We also remark that the scalars $E$ and $D$ are respectively associated with the intensity of the electric field and electric induction as perceived in the null frame, while $H$ and $B$ are their magnetic counterparts. In terms of the aligned tetrad invariants (8), the standard invariants take the form
@@
$\mathscr{P}+\textrm{i}\mathscr{Q}$ $\equiv\frac{1}{4}P_{\mu\nu}P^{\mu\nu}+\frac{\textrm{i}}{4}P_{\mu\nu}\star P^{\mu\nu}=-\frac{1}{2}(D+\textrm{i}H)^{2}.$ (9b)
@@
This highlights that in the more known Lagrangian formulation the fundamental variables are instead $E$ and $B$, being the others that must be determined from the constitutive relations. Using (10) as $\textrm{d}\mathscr{H}=-E\textrm{d}D+B\textrm{d}H$ in the differential of (12), the constitutive relations acquire now the alternative simple form
@@
which were first introduced in Ref. [31] and led to two alternative dual descriptions of nonlinear electrodynamics, using purely inductions or intensities as independent variables. As was first pointed out and exploited in [31], these dual formulations are precisely the ideal ones to transparently describe theories invariant under duality rotations, a fact that will be reviewed in Subsec III.1. Additionally, as was shown in [35] they are indispensable to determine the electrodynamics supporting the spinning nonlinearly charged black holes of [33, 34]. Correspondingly, the constitutive relations in the mixed representations are written as
@@
Regarding the energy-momentum tensor, its traceless part is formally written again as in (11b) but considering the new constitutive relations above, while according with the partial Legendre transforms (14) the trace changes to
@@
After clearly defining duality invariant theories through these formalisms in the next section, we shall apply the above expression to determine their zero trace subclass defining the ModMax theory.
@@ III A simpler derivation of the ModMax electrodynamics
In this section, we follow the formalism of the previous section to rederive the ModMax theory [30] in a novel and comprehensive way. Our approach will be based on the inductions and intensities formulations with structural functions $\mathscr{M}^{+}$ and $\mathscr{M}^{-}$, respectively, which in turn provide a direct and transparent way of deriving the theory from its defining symmetries. A symmetry-based approach was previously undertaken in [36] but employing a Lagrangian description. In contrast with the $\mathscr{M}^{\pm}$ formulations the resulting condition is nonlinear as we will explain later, and it allows no explicit general solution as recently emphasized in [37].
@@ III.1 Duality symmetry
whose general solutions are respectively of the form
@@
Hence, as was first realized in [31], a generic duality-invariant nonlinear electrodynamics is necessarily described by the above structural functions each depending in general on a single rotation-invariant variable built with the corresponding aligned invariants.
We end this subsection by a last comment regarding the duality condition $BE=DH$ in the Lagrangian and “Hamiltonian” formulations. Using the constitutive relations (13) or (10), this duality requisite can be expressed in either of the following two ways
@@
These are nonlinear first-order PDE of Hamilton-Jacobi type [38], and it is a nontrivial task to find their general solutions. In the case of the above equations, it is possible to find their general solutions depending on an arbitrary function, see Ref. [39] for a recent account of the several methods to achieve this task. Unfortunately, as occurs for all nonlinear first-order PDE, such general solutions are only implicitly given (see the Appendix and the recent Ref. [37]). This makes impossible a transparent treatment of duality invariant theories in the conventional formulations, i.e. the occurrence or not of duality invariance needs to be checked case by case.
@@ III.2 Conformal invariance
These linear PDE are easily solved and the result (which can also obtained in the other formulations) is that the structural functions characterizing all the conformally invariant nonlinear electrodynamics are necessarily homogeneous. For example, the degree of the previous functions $\mathscr{M}^{\pm}$ must be two, and by directly taking the trace of (5) it follows that the degree of $\mathscr{H}(\mathscr{P},\mathscr{Q})$ must be one. In order to exhibit that such conditions are enough to guarantee the conformal invariance of action (1), we explicitly write down the conformal weights of its different ingredients
@@
This together with the degree one homogeneity of the “Hamiltonian” imply the latter transforms with the appropriate weight, i.e.
@@ III.3 ModMax derivation
After having transparently characterized both duality and conformal symmetries, it will turn out to easily derive the ModMax theory [30]. The single-argument dependence of duality invariant theories (25) reduces the homogeneity defining PDE (27) warranting conformal symmetry to the simple ordinary equations
@@
These structural functions define the so-called ModMax electrodynamics, i.e. the generic family of duality invariant theories which enjoy at the same time conformal symmetry. The difference between the above integration constants is due to the fact that it is enough to integrate only one of the structural functions, since the other is obtained from the Legendre transforms (14) and results with the inverse integration constant of the former. In order to write the theory in a more standard form, notice that due to the squared dependence (31) the resulting constitutive relations (15) become linear and trivially invertible
@@
That turns into a straightforward task to obtain the other structural functions from the Legendre transforms (14) and (12) yielding also a squared dependence in terms of the aligned invariants
@@
We are now ready to express the Lagrangian and “Hamiltonian” in terms of the conventional invariants by inverting the parabolic relations (9) between the two families of invariants. This finally allows us to write down the standard structural functions for ModMax theory in their best known form [30]

### v2

@@ Abstract
We start this paper by concisely rederiving ModMax, which is nothing but the unique nonlinear extension of Maxwell’s equations preserving conformal and duality invariance. The merit of this new derivation is its transparency and simplicity since it is based on an approach where the elusive duality invariance is manifest. In the second part, we couple the ModMax electrodynamics to Einstein gravity with a cosmological constant together with a standard conformal scalar field, and new stationary spacetimes with dyonic charges are found. These solutions are later used as seed configurations to generate nonlinearly charged (super-)renormalizably dressed spacetimes by means of a known generating method that we extend to include any nonlinear conformal electrodynamics. We end by addressing the issue of how to generalize some of these results to include the recently studied non-Noetherian conformal scalar fields, whose equation of motion still enjoys conformal symmetry even though its action does not. It turns out that the static non-Noetherian conformally dressed black holes also become amenable to being charged by ModMax.
@@ I Introduction
It is certainly not an understatement to say that the paradigm of all effective low-energy field theories is due to the seminal works of Heisenberg, Euler, and Weisskopf [1, 2]. In more modern language, it is agreed that by integrating out the degrees of freedom of the electron field in quantum electrodynamics (QED), one obtains an approximate description of the photon field deviating from the standard linear regime [3]. This QED vacuum behavior makes nonlinear electrodynamics a semiphenomenological tool. As a result, several observational and experimental efforts have been deployed over the years to probe possible nonlinear effects in electrodynamics [4]. Well before these developments, Born and Infeld constructed a theory in which the electromagnetic field of a point charge is devoid of singularities [5]. Indeed, by going beyond the principle of superposition, they were able to solve the problem of infinite self-energy, and they argued that there should be a maximum field strength similar to the upper limit provided by the speed of light in Special Relativity. This reasoning was found to be consistent with string theory decades later, where Born-Infeld theory reappears in the dynamics of strings and branes, and this maximum field strength is precisely associated with the fact that such extended objects cannot move faster than the speed of light [6, 7]. This result further strengthens the interpretation of nonlinear electrodynamics as an effective field theory.
@@
The mentioned scale dependence is incompatible with conformal symmetry, one of the key features of standard Maxwell’s theory in four dimensions and in fact, of almost the whole standard model. Hence, it is also desirable to have explicit self-gravitating configurations involving other electrodynamics that depart from the Maxwell one but still preserve the conformal invariance. It is worth mentioning that static black holes are known for electromagnetism theories with conformal symmetry for dimensions $D\not=4$, which can be achieved by considering as Lagrangian a power law of the single field strength invariant with a dimension-dependent exponent [13]. These nonlinear models have also been shown to harmonize with conformally coupled scalar fields [14], even to the point of producing composite stealth configurations. [^1] By stealth we mean nontrivial field configurations with a vanishing energy-momentum tensor; see Refs. [15, 16, 17, 18, 19] It is natural to consider scalar fields at the same time since they are precisely the simplest way to realize conformal symmetry. In this respect, we mention that the most general second-order conformally invariant scalar field equation in four dimensions arising from an action principle was recently found in [20] after the precursor results of Ref. [21].
@@
In the context of black holes the relevance of conformally invariant fields has been evident since the pioneering and independent works of Refs. [22, 23], where the first example of a black hole dressed by a conformal scalar field was provided. This solution turns out to be pathological due to the divergence of the scalar field at the event horizon [24]. Remarkably, this pathology can be cured just by turning on a cosmological constant together with the unique conformally invariant potential in four dimensions [25]. This spherical black hole allows an electrically charged extension [25] and an analog whose event horizon is negatively curved [26]. For completeness we also mention that they have been generalized in several ways as solutions; see, e.g., [27, 28, 29, 30]. Of course, the dyonic extension of this so-called charged MTZ black hole is straightforward thanks to the duality and conformal symmetries of the standard Maxwell theory. Outstandingly, it turns out that Maxwell electrodynamics is not the only one that shares such properties. In fact, there exists a unique nonlinear extension of the Maxwell theory that inherited both its conformal and duality symmetry, which was recently proposed in Ref. [31] and dubbed ModMax. The main goal of this paper is to explore if the mere assumption of these two ingredients, i.e. duality as well as conformal invariance, is enough to find further generalizations of the known conformally dressed black holes. The results of our exploration are positive and, in what follows we summarize the different steps to achieve our objective.
@@
The plan of the paper is organized as follows. In order to be self-contained but also for usefulness, we start in Sec. II by providing a survey of nonlinear electrodynamics à la Salazar, García and Plebański [32]. This step is needed since it allows a transparent definition of duality-invariant theories that we will carefully justify in Sec. III.1. The conformal symmetry is then characterized in Sec. III.2 and later used beside the previous results to allow a clear rederivation of the ModMax theory in Sec. III.3. In Sec. IV, we provide the dyonic generalization, supplemented with a nontrivial NUT parameter, of the MTZ black hole nonlinearly charged through the ModMax theory. Since we go beyond Maxwell theory respecting power-counting renormalizability it makes total sense to do something similar with the scalar field. Fortunately, in Ref. [29] it was shown by means of a generating technique that the charged MTZ solution allows a generalization supported by a (super-)renormalizable self-interaction potential. Here, we first extend this generating technique to include any conformal nonlinear electrodynamics and then construct new (super-)renormalizably dressed spacetimes in Sec. V. This work is finally completed by adding to the scalar field a term that breaks its conformal invariance at the level of the action but not at the level of its equation of motion, i.e. we also address the study of the so-called non-Noetherian conformal scalar fields [21, 20]. In Sec. VI we succeed to charge with ModMax the static non-Noetherian conformally dressed black holes originally found by Fernandes in Ref. [21]. The last section, VII, is devoted to our conclusions.
@@ II Nonlinear electrodynamics
which depends on the metric $g_{\mu\nu}$, the gauge potential $A_{\mu}$ and the antisymmetric tensor $P^{\mu\nu}$. Here the structural function $\mathscr{H}$ describes the precise nonlinear electrodynamics and depends, in general, on the two Lorentz scalars that can be constructed with $P^{\mu\nu}$ [5]; see the first equality of Eq. (9b). As usual, the field strength is related to the gauge potential as $F=\textrm{d}A$, ensuring the Faraday equations
@@
with the subindices of $\mathscr{H}$ indicating partial differentiation, e.g., ${\mathscr{H}}_{\mathscr{P}}={\partial\mathscr{H}}/{\partial\mathscr{P}}$. Notice that Maxwell electrodynamics is recovered for $\mathscr{H}=\mathscr{P}$, giving linear constitutive relations. Lastly, the corresponding energy-momentum tensor reads
@@
The main motivation for the action principle (1) is that now Maxwell equations (3) remain linear as the Faraday ones (2) while the nonlinearity is encoded into the constitutive relations (4). Consequently, Maxwell equations (3) can be now understood just like the Faraday ones (2), i.e., implying the local existence of a vector potential $\star P=\textrm{d}A^{*}$. Therefore, from the point of view of the action principle (1), a solution to nonlinear electrodynamics can be understood as a pair of vector potentials $A$ and $A^{*}$ compatible with the constitutive relations (4). Additionally, since in four dimensions both Faraday (2) and Maxwell (3) equations define conservation laws, there are conserved quantities related to them defined by the following integrals
@@
After this brief and useful introduction, and in order to prepare for what follows, we review a strategy that has proved to be fruitful when nonlinear electrodynamics is considered in General Relativity [33]. This strategy simply consists of working in a null tetrad of the spacetime metric
@@
Here, the first pair of the tetrad is composed of complex conjugates one-forms, while the last pair is real. Additionally, it has been implicitly assumed that the electromagnetic configuration is algebraically general; namely, the real invariants $E$, $B$, $D$, and $H$ which are related to the eigenvalues are not all zero at the same time. We also remark that the scalars $E$ and $D$ are associated with the intensity of the electric field and electric induction, respectively, as perceived in the null frame, while $H$ and $B$ are their magnetic counterparts. In terms of the aligned tetrad invariants (8), the standard invariants take the form
@@
$\mathscr{P}+\textrm{i}\mathscr{Q}$ $\equiv\frac{1}{4}P_{\mu\nu}P^{\mu\nu}+\frac{\textrm{i}}{4}P_{\mu\nu}\star P^{\mu\nu}=-\frac{1}{2}(D+\textrm{i}H)^{2},$ (9b)
@@
This highlights that in the more known Lagrangian formulation the fundamental variables are instead $E$ and $B$, where the others must be determined from the constitutive relations. Using (10) as $\textrm{d}\mathscr{H}=-E\textrm{d}D+B\textrm{d}H$ in the differential of (12), the constitutive relations acquire now the alternative simple form
@@
which were first introduced in Ref. [32] and led to two alternative dual descriptions of nonlinear electrodynamics, using purely inductions or intensities as independent variables. As was first pointed out and exploited in [32], these dual formulations are precisely the ideal ones to transparently describe theories invariant under duality rotations, a fact that will be reviewed in Sec. III.1. Additionally, they are indispensable to determine the electrodynamics supporting the spinning nonlinearly charged black holes of [34, 35] as was shown in [36]. Correspondingly, the constitutive relations in the mixed representations are written as
@@
Regarding the energy-momentum tensor, its traceless part is formally written again as in (11b) but considering the new constitutive relations above, while the trace changes according to the partial Legendre transforms (14)
@@
After clearly defining duality-invariant theories through these formalisms in the next section, we shall apply the above expression to determine their zero trace subclass defining the ModMax theory.
@@ III A simpler derivation of the ModMax electrodynamics
In this section, we follow the formalism of the previous section to rederive the ModMax theory [31] in a novel and comprehensive way. Our approach will be based on the inductions and intensities formulations with structural functions $\mathscr{M}^{+}$ and $\mathscr{M}^{-}$, respectively, which in turn provide a direct and transparent way of deriving the theory from its defining symmetries. A symmetry-based approach was previously undertaken in [37] but employed a Lagrangian description. In contrast with the $\mathscr{M}^{\pm}$ formulations, the resulting condition is nonlinear as we will explain later, and it allows no explicit general solution as recently emphasized in [38].
@@ III.1 Duality symmetry
whose general solutions are, respectively, of the form
@@
Hence, as was first realized in [32], a generic duality-invariant nonlinear electrodynamics is necessarily described by the above structural functions, each generally depending on a single rotation-invariant variable built with the corresponding aligned invariants.
We end this subsection by a last comment regarding the duality condition $BE=DH$ in the Lagrangian and Hamiltonian formulations. Using the constitutive relations (13) or (10), this duality requisite can be expressed in either of the following two ways
@@
These are nonlinear first-order PDE of Hamilton-Jacobi type [39], and it is a nontrivial task to find their general solutions. In the case of the above equations, it is possible to find their general solutions depending on an arbitrary function; see Ref. [19] for a recent account of the several methods used to achieve this task. Unfortunately, as occurs for all nonlinear first-order PDE, such general solutions are only implicitly given (see the Appendix and the recent Ref. [38]). This makes impossible a transparent treatment of duality-invariant theories in the conventional formulations, i.e., the occurrence or not of duality invariance needs to be checked case by case.
@@ III.2 Conformal invariance
These linear PDE are easily solved and the result (which can also be obtained in the other formulations) is that the structural functions characterizing all the conformally invariant nonlinear electrodynamics are necessarily homogeneous. For example, the degree of the previous functions $\mathscr{M}^{\pm}$ must be two, and by directly taking the trace of (5) it follows that the degree of $\mathscr{H}(\mathscr{P},\mathscr{Q})$ must be one. In order to exhibit that such conditions are enough to guarantee the conformal invariance of action (1), we explicitly write down the conformal weights of its different ingredients
@@
They, together with the degree one homogeneity of the Hamiltonian, imply the latter transforms with the appropriate weight, i.e.,
@@ III.3 ModMax derivation
After having transparently characterized both duality and conformal symmetries, it will turn out to easily derive the ModMax theory [31]. The single-argument dependence of duality-invariant theories (25) reduces the homogeneity defining PDE (27), warranting conformal symmetry to the simple ordinary equations
@@
These structural functions define the so-called ModMax electrodynamics, i.e., the generic family of duality-invariant theories that enjoy conformal symmetry at the same time. The difference between the above integration constants is due to the fact that it is enough to integrate only one of the structural functions, since the other is obtained from the Legendre transforms (14) and results in the inverse integration constant of the former. In order to write the theory in a more standard form, notice that due to the squared dependence (31) the resulting constitutive relations (15) become linear and trivially invertible
@@
That turns into a straightforward task to obtain the other structural functions from the Legendre transforms (14) and (12), also yielding a squared dependence in terms of the aligned invariants
@@
We are now ready to express the Lagrangian and Hamiltonian in terms of the conventional invariants by inverting the parabolic relations (9) between the two families of invariants. This finally allows us to write down the standard structural functions for ModMax theory in their best known form [31]